OpenStudy (anonymous):

What is the simplest form of the fraction below? 4/12 a. 2/8 b. 2/3 c. 2/7 d. 1/3

OpenStudy (nurali):

\[\frac{ 4 }{ 12}\] \[4*3=12\] \[\frac{ 1 }{ 3 }\]